Make New Entry,
Make Followup Entry
User name Ole Hansen
Log entry time 15:57:56 on February 23, 2011
Entry number 348301
keyword=Day Shift Summary
Ole Hansen (SL)
Bo Zhao (TO/3rd)
Spent first 5 hours completing the configuration changes (spectrometer
polarity, angles, momenta) for the upcoming n-efficiency measurement.
Parasitically, various work in the hall was done. Second part of shift
consisted of data taking for the n-efficiency measurement (x>1, 500
- 08:00 Configuration change for n-efficiency in pogress
- 09:35-10:15 Problems with RHRS spectrometer motion. Rebooted IOC, reset
electronics. Eventually solved.
- 13:00 Access for configuration change completed. Requesting beam.
- 13:03-13:26 Raster check. LHRS BPMs don't seem to be working.Falling
back on spot_R
- 13:30 Begin data taking for n-efficiency check, runs 1769-1773
- 15:45 BB HV off for testing. Start run 1774.
- 16:00 Shift ends with run 1774 in progress.
- 1758-1766 cosmics, junk
- 1767-1768 spot++
- 1769 n-efficiency, 15 cm D2, 3 uA, 24 min, 280k T5
- 1770 n-efficiency, 15 cm D2, 3 uA, 25 min, 278k T5
- 1771 n-efficiency, 4 cm D2, 6 uA, 31 min, 119k T5
- 1772 n-efficiency, 4 cm D2, 6 uA, 31 min, 119k T5 (sic)
- 1773 n-efficiency, 4 cm D2, 6 uA, 15 min, 57k T5
- 1774 n-efficiency, 4 cm D2, 6 uA, BB HV off, in progress at end of